If a 15.6 g object requires 836.8 J of heat to raise its temperature by 21.0°C, what is the specific heat of the object? 


1
Expert's answer
2022-04-11T05:58:03-0400

q = m × C × ∆T

C = q / (m × ∆T) = 836.8 J / (15.6 g × 21.0°C) = 2.55 J/g °C